OpenStudy (anonymous):

I WILL AWARD MEDALS! Based on the graph below, what are the solutions to the equation f(x) = g(x)? graph of function f of x equals 4 times x minus 3. Graph of function g of x equals x cubed minus x squared minus 4 times x plus 4 a. x=-2.8, 0.6, -3.7 b. x=-2, 1, 2 c. x=-2, .5, 3.5 d. x=-2, 2, 3

OpenStudy (anonymous):

Here is the graph

OpenStudy (anonymous):

ok i'll look at it i'll figure it out

OpenStudy (anonymous):

okay :) thank you!

OpenStudy (anonymous):

@texaschic101 could you help me out?

OpenStudy (anonymous):

@AlohaKaika @pooja195

OpenStudy (anonymous):

I keep thinking it is A? I don't know if i'm right.

OpenStudy (anonymous):

Am I right? Is it A? @AlohaKaika

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

I think your right :)....hold on Ive gotta re-check though.

OpenStudy (anonymous):

Kk. <3

OpenStudy (anonymous):

Im pretty sure that your correct....I got A as well.....

OpenStudy (anonymous):

Okay! I was just getting a second opinion!

OpenStudy (anonymous):

Could you help me with a few more?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

Yay! Just give me one second.

OpenStudy (anonymous):

Emily wants to hang a painting in a gallery. The painting and frame must have an area of 31 square feet. The painting is 5 feet wide by 6 feet long. Which quadratic equation can be used to determine the thickness of the frame, x? Image of a frame 5 feet wide by 6 feet long, with an of x on the bottom and right sides a. 4x^2 + 22x - 1 = 0 b. 4x^2 + 22x + 31 = 0 c. x^2 + 11x - 1 = 0 d. x^2 + 11x + 31 = 0 Is it B?

OpenStudy (anonymous):

@AlohaKaika

OpenStudy (anonymous):

Uhhh...haha okay hold on :)

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

Hmm okay walk me through how you got B....just briefly....

OpenStudy (anonymous):

Well, i didn't know how to solve this problem. :/

OpenStudy (anonymous):

Ohhh gotcha....okay hold on...

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

You might have to ask someone else...haha :) Im sorry....I havent had to solve anything like that for quite a while.....sorry!!!

OpenStudy (anonymous):

its okay!

OpenStudy (anonymous):

I just don't want to give you the wrong anwser or anything!!!! :)

OpenStudy (anonymous):

ill figure it out! um, want to try a parabola problem i have trouble with?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

What is the domain of the following parabola? u-shaped graph that opens up with a vertex of (1, 0) y ≥ 0 All real numbers x ≥ 1 x ≤ 1

OpenStudy (anonymous):

Stupid question marks hold on

OpenStudy (anonymous):

y is greater than or equal to 0 All real numbers x is greater than or equal to 1 x is less than or equal to 1

OpenStudy (anonymous):

I was thinking it would be Choice B.

OpenStudy (anonymous):

Im pretty sure its B! :) Nice job....

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

Okay! <3

OpenStudy (anonymous):

One more before I go? :)

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

The steps to derive the quadratic formula are shown below: Step 1: ax2 + bx + c = 0 Step 2: x2 + b over ax + c over a = 0 Step 3: x2 + b over a x = - c over a Step 4: x2 +b over a x + square of quantity of b over 2 times a = -c over a + square of quantity of b over 2 times a Step 5: x plus b over 2 times a, all squared = -c over a + square of quantity of b over 2 times a Step 6: x + b over 2 times a = square root of all of negative c over a plus the square of the quantity b over 2 times a Step 7: x = - b over 2 times a ± square root of all of negative c over a plus the square of the quantity b over 4 times a = -b over 2 times a ± square root of the quantity of b squared minus 4 times a times c over 4 times a squared= negative b plus or minus square root of b squared minus 4 times a times c over 2 times a Which of the following statements best describes the first incorrect step? In Step 4, square of quantity of b over 2 times a instead of b over 2 times a was added on both sides. In Step 5, x plus b over 2 times a, all squared was written instead of x plus b over 4 times a, all squared . In Step 6, -c over a was written instead of c over a inside the root. In Step 7, b over 4 times a, all squared was written instead of square of quantity of b over 2 times a inside the root.

OpenStudy (anonymous):

HOLY COW!!! Haha thats alot of info.....

OpenStudy (anonymous):

Yeah I know. :P

OpenStudy (anonymous):

Geezee who gave you this problem? :)

OpenStudy (anonymous):

You can thank my teacher for that! :P its a worksheet

OpenStudy (anonymous):

I was thinking it would be A?

OpenStudy (anonymous):

Yup your right! :) because its supposed to be B over 2 :) so A is the anwser.

OpenStudy (anonymous):

YES! OMG THANK YOU! :D

OpenStudy (anonymous):

Haha sure thing! :)

OpenStudy (anonymous):

Okay, I'm gonna close this problem now! Here is a medal!

OpenStudy (anonymous):

Thanks! If you ever need anymore help let me know! :)

OpenStudy (anonymous):

Okay! <3